# Characterizing the Burst Error Correction Ability of Quantum Cyclic Codes

### v1

@@ I Introduction
Similar to classical coding theory [7], there are quantum burst error correction codes (QBECCs) [8, 9, 10, 11] for correcting bursts of errors. In [9], the spatially correlated qubit errors are considered and the first families of QBECCs were constructed by using the Calderbank-Shor-Steane (CSS) construction [12, 13]. In [11], QBECCs of short block length were found by using computer exhaustive search. In [10, 14], quantum interleavers for QBECCs were proposed so that long QBECCs can be produced from short ones. In [15], some QBECCs were obtained from quantum Reed-Solomon (RS) codes. In [16], the framework of general QBECCs was presented and degenerate QBECCs were first proposed. In [17, 18], quantum burst error correction was related to topological quantum error-correcting codes and new QBECCs were constructed by using interleaving techniques. More recently, burst errors caused by the cosmic rays are particularly concerned in the experiment of fault-tolerant quantum computations [19, 20, 21, 22, 23].
@@
Compared to the development of standard QECCs [24, 25, 26] or entanglement-assisted QECCs [27, 28, 29, 30, 31, 32, 33, 34, 35], the construction and investigation of QBECCs have received far less attention. The current QBECCs are mainly obtained with interleaving or by using computer exhaustive search. The interleaving technology highly relies on short QBECCs and the parameters are restricted to some specific numbers. The time complexity of exhaustively searching of general QBECCs are exponential to the code length [7]. In addition, there is an interesting class of quantum codes, called degenerate codes, that have no classical correspondences. Degenerate codes can potentially store more quantum information or correct more quantum errors than nondegenerate ones [1, 36, 37, 34]. However, determining the burst error correction ability of degenerate QBECCs is a quite difficult problem [16]. In particular, QBECCs saturating the quantum Reiger bound is the state-of-the-art in coding theory [7, 16]. On the other hand, almost all of quantum burst error correction codes are quantum cyclic codes (QCCs), which can be efficiently realized by using the quantum shift register [38]. This phenomenon learns from classical coding theory, in which cyclic codes are much more practical than non-cyclic codes [7]. How to determine the burst error correction limit of general QCCs is still unknown. Meanwhile, quantum RS codes are efficient to correct both random and burst errors. There exists the lower bound for the burst error correction ability of quantum RS codes [15, 7]. Whether the true burst error correction limit of quantum RS codes can exceed the lower bound is unknown.
@@
In this paper we characterize the burst error correction ability of QCCs by generalizing the algorithms in [39] for classical cyclic codes to QCCs. We propose polynomial-time algorithms to determine the burst error correction limit of general QCCs. Moreover, degenerate errors are particularly considered in the algorithms. As a result, a lot of new QBECCs which can achieve the quantum Reiger bound are obtained by running the polynomial-time searching algorithms in Magma software (V2.28-3) [40]. What’s more, we also propose polynomial-time algorithms to determine the true burst error correction limit of quantum RS codes. Many quantum RS codes with burst error correction ability beating the lower bound in [15, 7] are derived. At last, we propose the quantum error-trapping decoder (QETD) for correcting burst errors of QCCs. This decoding algorithm runs in linear-time and is quantum maximum likelihood decoding. In addition, the quantum error-trapping decoder can not only decode all the burst errors that are coset leaders but also can correct degenerate errors belong to the coset of coset leaders. Our numerical results show that QETD can decode much more degenerate burst errors than the nondegenerate ones. The Magma codes for computing the burst error correction limit of QCCs and quantum RS codes, and evaluating the performance of QETD are put in [41].
@@ II Preliminaries
[16] Let $C=[n,k]_{4}$ be an additive code over $\mathbb{F}_{4}$ and suppose that $C^{\bot_{H}}\subseteq C$, where $C^{\bot_{H}}$ is the Hermitian dual of $C$. Suppose that $\ell$ is the largest integer such that $e_{1}+e_{2}\notin C\backslash C^{\bot_{H}}$ for arbitrary two vectors $e_{1}\neq e_{2}\in\mathbb{F}_{4}^{n}$ with $0\leq\textrm{bl}(e_{1})$, $\textrm{bl}(e_{2})\leq\ell$. There exists a
$Q=[[n,2k-n]]$
@@ III The Burst Error Correction Ability of Quantum Cyclic Codes
We construct quantum cyclic codes from classical cyclic codes satisfying the dual containing relationship in Lemma 1 and Lemma 2. It should be noted that we mainly give the detailed process of determining the burst error correction ability of Hermitian-type QCCs. For other QCCs, e.g., the CSS-type QCCs, the process is simlilar to that of the Hermitian-type QCCs. Let ${C}=[n,k]_{q^{2}}$ be a classical cyclic code over $\mathbb{F}_{q^{2}}$. Denote $r=n-k$ by the number of check symbols. Denote $\mathbf{g}(x)=g_{0}+g_{1}x+\cdots+g_{r-1}x^{r-1}+g_{r}x^{r}$ and $\mathbf{h}(x)=h_{0}+h_{1}x+\cdots+h_{k-1}x^{k-1}+h_{k}x^{k}$ by the generator and the parity-check polynomials, respectively. The generator matrix of ${C}$ is given by
@@
where $11\leq i\leq r-$ 1. The parity-check matrix of ${C}$ is given by
@@
where $11\leq j\leq k-$ 1. Let $\mathcal{M}^{(t)}$ be the $(r-t)\times(n-t)$ matrix formed by deleting the last $t$ rows and last $t$ columns of the parity-check matrix $\mathbf{H}$ in (5), where $21\leq t\leq(n-k)/$ 2. Let $A=(a_{ij})$ be a matrix with elements over $\mathbb{F}_{q^{2}}$, where $1\leq i\leq m$ and $1\leq j\leq n$. The conjugate transpose of $A$ is given by $A^{\dagger}=(b_{uv}^{q})$, where $b_{uv}^{q}=a_{vu}^{q}$, $1\leq u\leq n$ and $1\leq v\leq m$. Then the conjugate transpose of $\mathbf{G}$ and $\mathbf{H}$ is denoted by $\mathbf{G}^{\dagger}$ and $\mathbf{H}^{\dagger}$, respectively. Firstly, we need the following three lemmas for determining the burst error correction ability of QCCs.
@@
[45, 49] Let $\textrm{gcd}(n,q^{2})=1$ and let ${C}=[n,k]_{q^{2}}$ be a classical cyclic code. If the parity-check matrix $\mathbf{H}$ of ${C}$ satisfying $\mathbf{H}\mathbf{H}^{\dagger}=0$, then ${C}^{\bot_{H}}\subseteq{C}$ and there exists a ${Q}=[[n,2k-n]]_{q}$ QCC.
@@
Let ${C}$ be a classical cyclic code over $\mathbb{F}_{q^{2}}$ with the parity-check matrix given in (5). Let $L$ be the largest integer $1\leq b\leq r$ such that every set of $b$ consecutive columns of the matrix $\mathcal{M}^{(b)}$ is linearly independent. Then ${C}$ can correct any burst error of length $L$ or less.
@@
through elementary row operations and a permutation of columns (if necessary), where $\mathbf{I}$ is an identity matrix of dimension $r$ equal to the rank of $M$, $X$ is a matrix with $r$ rows and $n-r$ columns, and the two $\mathbf{0}$’s are zero matrices.
For a nonzero matrix $M$ of size $m\times n$ and a set $A=[i,j]$ with $1\leq i\leq j\leq n$, we define the set of $j-i+1$ consecutive columns of $M$ indexed by $A$ as $M_{A}$. Denote the set of columns of $M$ by $\{m_{i}|1\leq i\leq n\}$, i.e., $M=(m_{i})_{1\leq i\leq n}$. We say that a set $\mathscr{S}$ of columns of $M$ is maximally linearly independent (MLI) if including any other column in $M$ would make it linearly dependent. If all the columns of $M$ is linearly independent, then $\mathscr{S}=\{m_{i}|1\leq i\leq n\}$. It is known that the rank of $M$ is equal to the number of elements in $\mathscr{S}$.
Let $1\leq\ell\leq r$ and let $A_{\imath}=[\imath,\imath+\ell-1]$, where $11\leq\imath\leq n-2\ell+$ 1. Denote a set of MLI columns in $\mathcal{M}^{(\ell)}_{A_{\imath}}$ by $\mathscr{D}_{H}=\{\alpha_{1},\cdots,\alpha_{u}\}$, where $\alpha_{u}\in\mathcal{M}^{(\ell)}_{A_{\imath}}$ and $1\leq u\leq\ell$ is the rank of $\mathcal{M}^{(\ell)}_{A_{\imath}}$. We define $\widehat{\mathscr{D}}_{H}=\{\beta_{1},\cdots,\beta_{v}\}$ as the complementary set of $\mathscr{D}_{H}$, where $v=\ell-u$. If $\mathcal{M}^{(\ell)}_{A_{\imath}}$ is of full rank, then $u=\ell$ and $\widehat{\mathscr{D}}_{H}$ is empty. Otherwise, each element in $\widehat{\mathscr{D}}_{H}$ can be represented by a linear combination of vectors in $\mathscr{D}_{H}$, i.e., $\beta_{i}=a_{i1}\alpha_{1}+\cdots+a_{iu}\alpha_{u}$ for $\beta_{i}\in\mathscr{D}_{H}$ and $1\leq i\leq v$. Denote by a burst error $e_{i}$ ($1\leq i\leq v$) whose nonzero components fall in the positions of $\{\alpha_{1},\cdots,\alpha_{u},\beta_{i}\}$ in $\mathbf{H}$. Then there exists a burst error $f_{i}$ whose nonzero components falling entirely in the last $\ell$ positions such that $e_{i}+f_{i}\in{C}$ for each $1\leq i\leq v$ according to Lemma 4. We define a set of $e_{i}$ and $f_{i}$ as
@@
[…]
00 $[[13,1]]$ $3$ $g=(\textbf{1}^{6}\textbf{2}^{5}\textbf{3}^{3}\textbf{2}^{1}\textbf{1}^{0})$
$[[15,3]]$ $3$ $g=(\textbf{1}^{6}\textbf{2}^{3}\textbf{1}^{0})$
$[[17,1]]$ $3$ $g=(\textbf{1}^{8}\textbf{3}^{7}\textbf{1}^{6}\textbf{1}^{5}\textbf{2}^{4}\textbf{1}^{3}\textbf{1}^{2}\textbf{3}^{1}\textbf{1}^{0})$
$[[25,5]]$ $5$ $g=(\textbf{1}^{10}\textbf{2}^{5}\textbf{1}^{0})$
$[[35,7]]$ $7$ $g=(\textbf{1}^{14}\textbf{3}^{7}\textbf{1}^{0})$
$[[39,3]]$ $9$ $g=(\textbf{1}^{18}\textbf{3}^{15}\textbf{2}^{9}\textbf{3}^{3}\textbf{1}^{0})$
$[[45,9]]$ $9$ $g=(\textbf{1}^{18}\textbf{2}^{9}\textbf{1}^{0})$
$[[51,19]]$ $8$ $g=(\textbf{1}^{16}\textbf{2}^{14}\textbf{2}^{12}\textbf{1}^{8}\textbf{2}^{6}\textbf{2}^{5}\textbf{3}^{4}\textbf{3}^{3}\textbf{3}^{2}\textbf{3}^{1}\textbf{1}^{0})$
@@ 00 $[[13,1]]$ $3$ $g=(\textbf{1}^{6}\textbf{2}^{5}\textbf{3}^{3}\textbf{2}^{1}\textbf{1}^{0})$
$[[63,27]]$ $9$ $g=(\textbf{1}^{18}\textbf{3}^{17}\textbf{2}^{16}\textbf{1}^{15}\textbf{2}^{14}\textbf{2}^{13}\textbf{3}^{12}\textbf{1}^{11}\textbf{2}^{10}\textbf{3}^{9}\textbf{1}^{6}\textbf{1}^{5}\textbf{1}^{3}\textbf{3}^{1}\textbf{3}^{0}$
@@
$[[65,29]]$ $9$ $g=(\textbf{1}^{18}\textbf{3}^{17}\textbf{1}^{15}\textbf{3}^{13}\textbf{3}^{12}\textbf{2}^{11}\textbf{3}^{9}\textbf{2}^{7}\textbf{3}^{6}\textbf{3}^{5}\textbf{1}^{3}\textbf{3}^{1}\textbf{1}^{0})$
@@
$[[85,9]]$ $19$ $g=(\textbf{1}^{38}\textbf{1}^{36}\textbf{3}^{35}\textbf{3}^{34}\textbf{1}^{33}\textbf{1}^{31}\textbf{1}^{29}\textbf{2}^{28}\textbf{3}^{27}\textbf{3}^{26}\textbf{3}^{24}\textbf{1}^{23}\textbf{3}^{22}\textbf{3}^{21}\textbf{1}^{20}\textbf{3}^{19}\textbf{1}^{18}\textbf{1}^{17}\textbf{1}^{16}\textbf{3}^{14}\textbf{3}^{13}\textbf{2}^{10}\textbf{3}^{8}\textbf{1}^{6}\textbf{2}^{5}\textbf{1}^{4}\textbf{2}^{3}\textbf{2}^{2}\textbf{1}^{1}\textbf{1}^{0})$
@@
$[[85,49]]$ $9$ $g=(\textbf{1}^{18}\textbf{3}^{17}\textbf{2}^{16}\textbf{1}^{15}\textbf{1}^{14}\textbf{3}^{13}\textbf{3}^{12}\textbf{1}^{11}\textbf{2}^{10}\textbf{1}^{9}\textbf{2}^{8}\textbf{3}^{7}\textbf{3}^{6}\textbf{3}^{5}\textbf{2}^{4}\textbf{1}^{3}\textbf{3}^{1}\textbf{1}^{0})$
$[[85,53]]$ $8$ $g=(\textbf{1}^{16}\textbf{3}^{15}\textbf{2}^{14}\textbf{1}^{11}\textbf{1}^{10}\textbf{3}^{9}\textbf{1}^{8}\textbf{1}^{7}\textbf{2}^{6}\textbf{3}^{5}\textbf{3}^{4}\textbf{3}^{2}\textbf{1}^{0})$
$[[91,7]]$ $21$ $g=(\textbf{1}^{42}\textbf{3}^{35}\textbf{2}^{21}\textbf{3}^{7}\textbf{1}^{0})$
$[[93,13]]$ $20$ $g_{1}=(\textbf{1}^{6}\textbf{1}^{4}\textbf{1}^{1}\textbf{1}^{0})$, $g_{2}=(\textbf{1}^{6}\textbf{1}^{4}\textbf{1}^{2}\textbf{1}^{1}\textbf{1}^{0})$

### v2

@@ I Introduction
Similar to classical coding theory [37], there are quantum burst error correction codes (QBECCs) [19, 54, 31, 49] for correcting bursts of errors. In [54], the spatially correlated qubit errors are considered and the first families of QBECCs were constructed by using the Calderbank-Shor-Steane (CSS) construction [6, 47]. In [49], QBECCs of short block length were found by using computer exhaustive search. In [31, 50], quantum interleavers for QBECCs were proposed so that long QBECCs can be produced from short ones. In [45], some QBECCs were obtained from quantum Reed-Solomon (RS) codes. In [15], the framework of general QBECCs was presented and degenerate QBECCs were first proposed. In [51, 52], quantum burst error correction was related to topological quantum error-correcting codes and new QBECCs were constructed by using interleaving techniques. In [12], a new class of quantum burst error-locating codes was proposed to locate and correct burst errors.
@@
Compared to the development of standard QECCs [43, 36, 13] or entanglement-assisted QECCs [34, 3, 21, 55, 28, 25, 26, 18, 14], the construction and investigation of QBECCs have received far less attention. Although standard QECCs can also be used to correct burst errors, they are not efficient enough and usually have a much smaller code rate than QBECCs for correcting burst errors of the same length. The current QBECCs are mainly obtained with interleaving or by using computer exhaustive search. The interleaving technology highly relies on short QBECCs and the parameters are restricted to some specific numbers. The time complexity of exhaustively searching of general QBECCs are exponential to the code length [37]. In addition, there is an interesting class of quantum codes, called degenerate codes, that have no classical correspondences. Degenerate codes can potentially store more quantum information or correct more quantum errors than nondegenerate ones [41, 17, 16, 18]. However, determining the burst error correction ability of degenerate QBECCs is a quite difficult problem [15]. In particular, QBECCs saturating the quantum Reiger bound is the state-of-the-art in coding theory [37, 15]. On the other hand, almost all of quantum burst error correction codes are quantum cyclic codes (QCCs), which can be efficiently realized by using the quantum shift register [23]. This phenomenon learns from classical coding theory, in which cyclic codes are much more practical than non-cyclic codes [37]. How to determine the burst error correction limit of general QCCs is still unknown. Meanwhile, quantum RS codes are efficient to correct both random and burst errors. There exists the lower bound for the burst error correction ability of quantum RS codes [45, 37]. Whether the true burst error correction limit of quantum RS codes can exceed the lower bound is unknown.
@@
In this paper we characterize the burst error correction ability of QCCs by generalizing the algorithms in [39] for classical cyclic codes to QCCs. We propose polynomial-time algorithms to determine the burst error correction limit of general QCCs. Moreover, degenerate errors are particularly considered in the algorithms. As a result, a lot of new QBECCs which can achieve the quantum Reiger bound are obtained by running the polynomial-time searching algorithms in Magma software (V2.28-3) [7]. What’s more, we also propose polynomial-time algorithms to determine the true burst error correction limit of quantum RS codes. Many quantum RS codes with burst error correction ability beating the lower bound in [45, 37] are derived. At last, we propose the quantum error-trapping decoder (QETD) for correcting burst errors of QCCs. This decoding algorithm runs in linear-time and is quantum maximum likelihood decoding. In addition, the quantum error-trapping decoder can not only decode all the burst errors that are coset leaders but also can correct degenerate errors which belong to the coset of coset leaders. Our numerical results show that QETD can decode much more degenerate burst errors than the nondegenerate ones. The Magma codes for computing the burst error correction limit of QCCs and quantum RS codes, and evaluating the performance of QETD are put in [20].
@@ II Preliminaries
[…]
Q=[[n,2kn]]Q=[[n,2k-n]]
@@ III The Burst Error Correction Ability of Quantum Cyclic Codes
We construct quantum cyclic codes from classical cyclic codes satisfying the dual containing relationship in Lemma 1 and Lemma 2. It should be noted that we mainly give the detailed process of determining the burst error correction ability of Hermitian-type QCCs. For other QCCs, e.g., the CSS-type QCCs, the process is simlilar to that of the Hermitian-type QCCs. Let ${C}=[n,k]_{q^{2}}$ be a classical cyclic code over $\mathbb{F}_{q^{2}}$. Denote r=nkr=n-k by the number of check symbols. Denote $\mathbf{g}(x)=g_{0}+g_{1}x+\cdots+g_{r-1}x^{r-1}+g_{r}x^{r}$ and $\mathbf{h}(x)=h_{0}+h_{1}x+\cdots+h_{k-1}x^{k-1}+h_{k}x^{k}$ by the generator and the parity-check polynomials, respectively. The generator matrix of C{C} is given by
@@
where $1\leq i\leq r-1$. The parity-check matrix of C{C} is given by
@@
where $1\leq j\leq k-1$. Let $\mathcal{M}^{(t)}$ be the $(r-t)\times(n-t)$ matrix formed by deleting the last $t$ rows and last $t$ columns of the parity-check matrix $\mathbf{H}$ in (5), where $1\leq t\leq(n-k)/2$. Let $A=(a_{ij})$ be a matrix with elements over $\mathbb{F}_{q^{2}}$, where $1\leq i\leq m$ and $1\leq j\leq n$. The conjugate transpose of $A$ is given by $A^{\dagger}=(b_{uv}^{q})$, where $b_{uv}^{q}=a_{vu}^{q}$, $1\leq u\leq n$ and $1\leq v\leq m$. Then the conjugate transpose of $\mathbf{G}$ and $\mathbf{H}$ is denoted by $\mathbf{G}^{\dagger}$ and $\mathbf{H}^{\dagger}$, respectively. Firstly, we need the following three lemmas for determining the burst error correction ability of QCCs.
@@
[5, 32] Let $\textrm{gcd}(n,q^{2})=1$ and let ${C}=[n,k]_{q^{2}}$ be a classical cyclic code. If the parity-check matrix $\mathbf{H}$ of C{C} satisfying $\mathbf{H}\mathbf{H}^{\dagger}=0$, then ${C}^{\bot_{H}}\subseteq{C}$ and there exists a ${Q}=[[n,2k-n]]_{q}$ QCC.
@@
Let C{C} be a classical cyclic code over $\mathbb{F}_{q^{2}}$ with the parity-check matrix given in (5). Let $L$ be the largest integer $1\leq b\leq r$ such that every set of $b$ consecutive columns of the matrix $\mathcal{M}^{(b)}$ is linearly independent. Then C{C} can correct any burst error of length $L$ or less.
@@
through elementary row operations and a permutation of columns (if necessary), where $\mathbf{I}$ is an identity matrix of dimension $r$ equal to the rank of $M$, $X$ is a matrix with $r$ rows and nrn-r columns, and the two $\mathbf{0}$’s are zero matrices.
For a nonzero matrix $M$ of size $m\times n$ and a set $A=[i,j]$ with $1\leq i\leq j\leq n$, we define the set of ji+1j-i+1 consecutive columns of $M$ indexed by $A$ as $M_{A}$. Denote the set of columns of $M$ by $\{m_{i}|1\leq i\leq n\}$, i.e., $M=(m_{i})_{1\leq i\leq n}$. We say that a set $\mathscr{S}$ of columns of $M$ is maximally linearly independent (MLI) if including any other column in $M$ would make it linearly dependent. If all the columns of $M$ is linearly independent, then $\mathscr{S}=\{m_{i}|1\leq i\leq n\}$. It is known that the rank of $M$ is equal to the number of elements in $\mathscr{S}$.
Let $1\leq\ell\leq r$ and let $A_{\imath}=[\imath,\imath+\ell-1]$, where $1\leq\imath\leq n-2\ell+1$. Denote a set of MLI columns in $\mathcal{M}^{(\ell)}_{A_{\imath}}$ by $\mathscr{D}_{H}=\{\alpha_{1},\cdots,\alpha_{u}\}$, where $\alpha_{u}\in\mathcal{M}^{(\ell)}_{A_{\imath}}$ and $1\leq u\leq\ell$ is the rank of $\mathcal{M}^{(\ell)}_{A_{\imath}}$. We define $\widehat{\mathscr{D}}_{H}=\{\beta_{1},\cdots,\beta_{v}\}$ as the complementary set of $\mathscr{D}_{H}$, where $v=\ell-u$. If $\mathcal{M}^{(\ell)}_{A_{\imath}}$ is of full rank, then $u=\ell$ and $\widehat{\mathscr{D}}_{H}$ is empty. Otherwise, each element in $\widehat{\mathscr{D}}_{H}$ can be represented by a linear combination of vectors in $\mathscr{D}_{H}$, i.e., $\beta_{i}=a_{i1}\alpha_{1}+\cdots+a_{iu}\alpha_{u}$ for $\beta_{i}\in\mathscr{D}_{H}$ and $1\leq i\leq v$. Denote by a burst error $e_{i}$ ($1\leq i\leq v$) whose nonzero components fall in the positions of $\{\alpha_{1},\cdots,\alpha_{u},\beta_{i}\}$ in $\mathbf{H}$. Then there exists a burst error $f_{i}$ whose nonzero components falling entirely in the last $\ell$ positions such that $e_{i}+f_{i}\in{C}$ for each $1\leq i\leq v$ according to Lemma 4. We define a set of $e_{i}$ and $f_{i}$ as
@@
$\Delta$ $[[n,k]]$ $\mathcal{L}$ Generator Polynomials
0 $[[13,1]]$ 33 $g=(\textbf{1}^{6}\textbf{2}^{5}\textbf{3}^{3}\textbf{2}^{1}\textbf{1}^{0})$
$[[15,3]]$ 33 $g=(\textbf{1}^{6}\textbf{2}^{3}\textbf{1}^{0})$
$[[17,1]]$ 33 $g=(\textbf{1}^{8}\textbf{3}^{7}\textbf{1}^{6}\textbf{1}^{5}\textbf{2}^{4}\textbf{1}^{3}\textbf{1}^{2}\textbf{3}^{1}\textbf{1}^{0})$
$[[25,5]]$ 55 $g=(\textbf{1}^{10}\textbf{2}^{5}\textbf{1}^{0})$
$[[35,7]]$ 77 $g=(\textbf{1}^{14}\textbf{3}^{7}\textbf{1}^{0})$
$[[39,3]]$ 99 $g=(\textbf{1}^{18}\textbf{3}^{15}\textbf{2}^{9}\textbf{3}^{3}\textbf{1}^{0})$
$[[45,9]]$ 99 $g=(\textbf{1}^{18}\textbf{2}^{9}\textbf{1}^{0})$
$[[51,19]]$ 88 $g=(\textbf{1}^{16}\textbf{2}^{14}\textbf{2}^{12}\textbf{1}^{8}\textbf{2}^{6}\textbf{2}^{5}\textbf{3}^{4}\textbf{3}^{3}\textbf{3}^{2}\textbf{3}^{1}\textbf{1}^{0})$
@@ 0 $[[13,1]]$ 33 $g=(\textbf{1}^{6}\textbf{2}^{5}\textbf{3}^{3}\textbf{2}^{1}\textbf{1}^{0})$
$[[63,27]]$ 99 $g=(\textbf{1}^{18}\textbf{3}^{17}\textbf{2}^{16}\textbf{1}^{15}\textbf{2}^{14}\textbf{2}^{13}\textbf{3}^{12}\textbf{1}^{11}\textbf{2}^{10}\textbf{3}^{9}\textbf{1}^{6}\textbf{1}^{5}\textbf{1}^{3}\textbf{3}^{1}\textbf{3}^{0}$
@@
$[[65,29]]$ 99 $g=(\textbf{1}^{18}\textbf{3}^{17}\textbf{1}^{15}\textbf{3}^{13}\textbf{3}^{12}\textbf{2}^{11}\textbf{3}^{9}\textbf{2}^{7}\textbf{3}^{6}\textbf{3}^{5}\textbf{1}^{3}\textbf{3}^{1}\textbf{1}^{0})$
@@
$[[85,9]]$ 1919 $g=(\textbf{1}^{38}\textbf{1}^{36}\textbf{3}^{35}\textbf{3}^{34}\textbf{1}^{33}\textbf{1}^{31}\textbf{1}^{29}\textbf{2}^{28}\textbf{3}^{27}\textbf{3}^{26}\textbf{3}^{24}\textbf{1}^{23}\textbf{3}^{22}\textbf{3}^{21}\textbf{1}^{20}\textbf{3}^{19}\textbf{1}^{18}\textbf{1}^{17}\textbf{1}^{16}\textbf{3}^{14}\textbf{3}^{13}\textbf{2}^{10}\textbf{3}^{8}\textbf{1}^{6}\textbf{2}^{5}\textbf{1}^{4}\textbf{2}^{3}\textbf{2}^{2}\textbf{1}^{1}\textbf{1}^{0})$
@@
$[[85,49]]$ 99 $g=(\textbf{1}^{18}\textbf{3}^{17}\textbf{2}^{16}\textbf{1}^{15}\textbf{1}^{14}\textbf{3}^{13}\textbf{3}^{12}\textbf{1}^{11}\textbf{2}^{10}\textbf{1}^{9}\textbf{2}^{8}\textbf{3}^{7}\textbf{3}^{6}\textbf{3}^{5}\textbf{2}^{4}\textbf{1}^{3}\textbf{3}^{1}\textbf{1}^{0})$
$[[85,53]]$ 88 $g=(\textbf{1}^{16}\textbf{3}^{15}\textbf{2}^{14}\textbf{1}^{11}\textbf{1}^{10}\textbf{3}^{9}\textbf{1}^{8}\textbf{1}^{7}\textbf{2}^{6}\textbf{3}^{5}\textbf{3}^{4}\textbf{3}^{2}\textbf{1}^{0})$
$[[91,7]]$ 2121 $g=(\textbf{1}^{42}\textbf{3}^{35}\textbf{2}^{21}\textbf{3}^{7}\textbf{1}^{0})$
$[[93,13]]$ 2020 $g_{1}=(\textbf{1}^{6}\textbf{1}^{4}\textbf{1}^{1}\textbf{1}^{0})$, $g_{2}=(\textbf{1}^{6}\textbf{1}^{4}\textbf{1}^{2}\textbf{1}^{1}\textbf{1}^{0})$